OpenStudy (anonymous):

Using variable substitution, how do I find the expression for y^2 - 2y + 3, if y = 4 + h

OpenStudy (amistre64):

replace y with (4+h) and expand

OpenStudy (amistre64):

you should get a quadratic in terms of h, then test the solutions to see if you have to abandon one of them

OpenStudy (anonymous):

(4+h)^2 - 2(4 + h) + 3

OpenStudy (amistre64):

thats the first step yes

OpenStudy (anonymous):

I got 16 + h^2 - 5 - 2h

OpenStudy (amistre64):

we could prolly use that as is :) \[4+h = \frac{2\pm\sqrt{4-4(3)}}{2}\]

OpenStudy (anonymous):

So that is the final answer? Is my answer right though?

OpenStudy (amistre64):

(4+h)^2 - 2(4 + h) + 3 16 +8h +h^2 -8 -2h + 3 11 +6h +h^2

OpenStudy (amistre64):

\[h=\frac{-6\pm\sqrt{36-4(11)}}{2}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

but the expression is the expansion itself

OpenStudy (anonymous):

So if I wanted to do a question like 3m^2 - 7, if m = 1 - k, what would I do?

OpenStudy (amistre64):

same thing, replace ms by (1-k) and expand it out

OpenStudy (anonymous):

so it would be 3(1 - k) (1 - k) - 7

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

So what would the result be if I expand that? Sorry, I am bad with this stuff!

OpenStudy (amistre64):

give it a try, since practice makes perfect

OpenStudy (anonymous):

I keep on forgetting what to do with the 3

OpenStudy (amistre64):

either use it first in the multiplication process, or use it last ...

OpenStudy (amistre64):

abc = (ab)c = a(bc) = (ac)b

OpenStudy (anonymous):

I don't really get it

OpenStudy (anonymous):

It's okay. Got it

OpenStudy (anonymous):

The answer is 3k^2 - 6k - 4

OpenStudy (amistre64):

3*(4*6) = (3*4)*6 yes, that looks correct

OpenStudy (anonymous):

:) Thanks!

OpenStudy (amistre64):

good luck